For the velocity potential function, φ = x^2 - y^2 . Find the velocity components at the point (4, 5).
1
Expert's answer
2020-12-14T09:35:50-0500

In velocity we can write 


"u=\dfrac{\partial\varphi}{\partial x}, v=\frac{\partial\varphi}{\partial y}"


Given "\varphi=x^2-y^2"

"u=\dfrac{\partial\varphi}{\partial x}=2x, v=\dfrac{\partial\varphi}{\partial y}=-2y"

Calculate the velocity at the point (4, 5)


"u=8, v=-10"

"|velocity|=\sqrt{u^2+v^2 }=\sqrt{(8)^2+(-10)^2}=2\sqrt{41}"